A certain mountain has an elevation of 19,785ft. In 1919, the glacier on this peak covered 10 acres. By 2003, this glacier has melted to 1 acre. The yearly rate of change is ___ acres/year. What is the equation that gives the acreage of the glacier (t) years after 1919? Any help would be appeaciated. Thanks
Answer by stanbon(75887)

A certain mountain has an elevation of 19,785ft. In 1919, the glacier on this peak covered 10 acres. By 2003, this glacier has melted to 1 acre. The yearly rate of change is ___ acres/year. What is the equation that gives the acreage of the glacier (t) years after 1919? Any help would be appeaciated.
----
You have 2 points relating year after 1919 and acreage:
(0,10) and (3,1)
-----
slope = (1-10)/(3-0) = -3
---
intercept = G(0) = 10
----------------------------
Equation:
G(t) = -3t + 10
=====================
Cheers,
Stan H.
=====================